7. Express sin 67° + cos 75° in terms of trigonometric ratios of angles between 0° and 45°.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given: sin 67° + cos 75° In term of sin as cos function and cos as sin function, it can be written as follows sin 67° = sin (90° – 23°) cos 75° = cos (90° – 15°) So, sin 67° + cos 75° = sin (90° – 23°) + cos (90° – 15°) Now, simplify the above equation = cos 23° + sin 15° Therefore, sin 67° + cos 75° is also expressed as cos 23° + sin 15°